# Approximate independence of permutation mixtures

### v2

@@ Appendix C Deferred proofs
=$\overset{\rm(d)}{=}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\binom{n-\ell}{k-\ell}^{2}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\frac{(k-\ell)!}{n^{k-\ell}},$
@@
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})+1$ $=\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})$
=$\overset{\rm(e)}{=}\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\cdot\frac{n!}{(n-\ell)!n^{\ell}}S_{\ell}=\sum_{\ell=0}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell},$
where (e) follows from the identity between $T_{\ell}$ and $S_{\ell}$ in Lemma A.1. Finally, as $S_{0}=1$, $S_{1}=0$, and $S_{\ell}\geq 0$,
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})=\sum_{\ell=2}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell}\leq\frac{k(k-1)}{n(n-1)}\sum_{\ell=2}^{n}S_{\ell}\overset{\rm(f)}{=}\frac{k(k-1)}{n(n-1)}\chi^{2}(P_{Y^{n}}\|Q_{Y^{n}}),$
@@
where (f) is due to the identity (27). Now the claimed result follows from the upper bounds of $\chi^{2}(P_{Y^{n}}\|Q_{Y^{n}})$ in Theorem 1.2.
C.4 Proof of Corollary 1.7
We first show that
@@
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right].$ (45)
@@
$r^{\mathrm{S}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{S}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\Delta(\cdot)}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\Delta(X_{i}))\right]=\mathbb{E}\left[\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]\right],$
$r^{\mathrm{PI}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\widehat{\vartheta}_{i}(X))\right]=\mathbb{E}\left[\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right],$
@@
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ $\leq\sum_{i=1}^{n}\mathbb{E}\left[\sup_{u}\left|\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]-\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right|\right]$
$\leq M\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right],$
@@
$\frac{\mathrm{d}P_{\vartheta_{i}|X_{i}}}{\mathrm{d}P_{\vartheta_{i}|X}}=\frac{\mathrm{d}P_{X_{\backslash i}|X_{i}}}{\mathrm{d}P_{X_{\backslash i}|\vartheta_{i}}},$
@@
$\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right]$ $=\mathbb{E}_{\vartheta_{i},X}\left[\left|\frac{\mathrm{d}P_{X_{\backslash i}|X_{i}}}{\mathrm{d}P_{X_{\backslash i}|\vartheta_{i}}}-1\right|\right]=\mathbb{E}_{\vartheta_{i},X_{i}}\left[{\sf TV}(P_{X_{\backslash i}|X_{i}},P_{X_{\backslash i}|\vartheta_{i}})\right]$
$\overset{\rm(a)}{\leq}\mathbb{E}_{X_{i}}\mathbb{E}_{\vartheta_{i},\vartheta_{i}^{\prime}|X_{i}}\left[{\sf TV}(P_{X_{\backslash i}|\vartheta_{i}^{\prime}},P_{X_{\backslash i}|\vartheta_{i}})\right]\leq\max_{\vartheta_{i},\vartheta_{i}^{\prime}}{\sf TV}(P_{X_{\backslash i}|\vartheta_{i}},P_{X_{\backslash i}|\vartheta_{i}^{\prime}})$

### v3

@@ Appendix C Deferred proofs
$\overset{\rm(d)}{=}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\binom{n-\ell}{k-\ell}^{2}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\frac{(k-\ell)!}{n^{k-\ell}},$
@@
$\chi^{2}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n})+1$ $=\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})$
$\overset{\rm(e)}{=}\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\cdot\frac{n!}{(n-\ell)!n^{\ell}}S_{\ell}=\sum_{\ell=0}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell},$
where (e) follows from the identity between $T_{\ell}$ and $S_{\ell}$ in Lemma˜A.1.
Finally, to prove the first upper bound, by (29) we have
$S_{\ell}$ $=\frac{n^{n}}{n!}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\frac{\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})^{2}}{(n-\ell)!\ell_{2}!\cdots\ell_{n}!}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}$
$\overset{\rm(f)}{\leq}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\frac{10\cdot\ell!(n-\ell)!}{(n-\ell)!\ell_{2}!\cdots\ell_{n}!}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}=10\left(\sum_{i=2}^{n}\lambda_{i}\right)^{\ell}\leq 10{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell},$
where (f) follows from the same program (Lemma˜4.2 and Theorem˜4.3) in Section˜4, and the last inequalty uses Lemma˜5.2. As $S_{0}=1$, $S_{1}=0$,
$\chi^{2}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n})\leq\sum_{\ell=2}^{k}\left(\frac{k}{n}\right)^{\ell}\cdot 10{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}\leq 20\left(\frac{k{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}{n}\right)^{2},\quad\text{if }\frac{k{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}{n}\leq\frac{1}{2}.$
For the second upper bound, we use (29) and $S_{\ell}\geq 0$ to conclude
$\chi^{2}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k.n})=\sum_{\ell=2}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell}\leq\frac{k(k-1)}{n(n-1)}\sum_{\ell=2}^{n}S_{\ell}=\frac{k(k-1)}{n(n-1)}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n}).$
@@
Now the claimed result follows from the upper bound of $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ in Theorem˜1.2.
C.4 Proof of Theorem˜1.5
By Lemma˜5.5, we have
where (a) is due to $\prod_{i=2}^{n-1}(1+\ell_{i})\geq 1+\sum_{i=2}^{n-1}\ell_{i}=\ell$, (b) uses $2\sqrt{\ell+1}\leq\ell+2$ for all $\ell\geq 0$, and (c) follows from the identity
By the same arguments in Section˜C.1, Lemma˜5.6 leads to
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})^{2}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})},$
which establishes the first claim of Theorem˜1.5. The second claim of Theorem˜1.5 simply follows from the Cauchy–Schwarz inequality.
C.5 Proof of Lemma˜2.1
Let $\mathbb{Q}$ be the joint distribution of $(\theta,X)$ under the i.i.d. prior $\theta\sim\left((1-\alpha)\delta_{0}+\alpha\delta_{\mu_{0}}\right)^{\otimes n}$, and $\mathbb{P}$ be the counterpart under the permutation prior. By the arguments in [Zha12, Proof of Theorem 1], to establish the claimed result, it suffices to prove that
$\mathbb{E}_{\mathbb{P}}\left[\mathbb{E}_{\mathbb{P}}[\theta_{1}|X^{n}]^{2}\right]=o\left(\frac{s\mu^{2}}{n}\right).$ (48)
For completeness we show how (48) implies the target Bayes risk lower bound. Let $\mu_{0}:=\sqrt{1-\varepsilon}\mu$. For $p_{1}\in(0,1)$, consider the minimizer $a^{\star}$ of the map $a\mapsto f(a):=p_{1}|\mu_{0}-a|^{q}+(1-p_{1})|a|^{q}$. Clearly $a^{\star}\in[0,\mu_{0}]$, and since $(1-p_{1})(a^{\star})^{q}\leq f(a^{\star})\leq f(0)=p_{1}\mu_{0}^{q}$, we conclude that $a^{\star}\leq(\frac{p_{1}}{1-p_{1}})^{1/q}\mu_{0}$. Consequently, if $p_{1}\leq\varepsilon$, then
$f(a^{\star})\geq p_{1}(1-\varepsilon_{1})^{q}\mu_{0}^{q},\quad\text{with }\varepsilon_{1}:=\left(\frac{\varepsilon}{1-\varepsilon}\right)^{1/q}.$
Therefore, letting $p_{1}(X^{n}):=\mathbb{P}(\theta_{1}=\mu_{0}|X^{n})$ be the posterior probability of $\theta_{1}=\mu_{0}$, the Bayes $\ell^{q}$ risk for estimating $\theta_{1}$ under $\mathbb{P}$ is
$\mathbb{E}_{\mathbb{P}}\left[\min_{a}\left(p_{1}(X^{n})|\mu-a|^{q}+(1-p_{1}(X^{n}))|a|^{q}\right)\right]\geq(1-\varepsilon_{1})^{q}\mu_{0}^{q}\cdot\mathbb{E}_{\mathbb{P}}\left[p_{1}(X^{n})\mathbbm{1}(p_{1}(X^{n})\leq\varepsilon)\right].$
To proceed, note that $\mathbb{E}_{\mathbb{P}}[p_{1}(X^{n})]=\mathbb{P}(\theta_{1}=\mu_{0})=\frac{\lfloor s\rfloor}{n}$, and $\mathbb{E}_{\mathbb{P}}[p_{1}(X^{n})^{2}]=o(\frac{s}{n})$ by (48). Therefore, the Bayes $\ell^{q}$ risk for estimating $\theta_{1}$ is further lower bounded by
$(1-\varepsilon_{1})^{q}\mu_{0}^{q}\cdot\left(\frac{\lfloor s\rfloor}{n}-\mathbb{E}_{\mathbb{P}}[p_{1}(X^{n})\mathbbm{1}(p_{1}(X^{n})>\varepsilon)]\right)$
$\geq(1-\varepsilon_{1})^{q}\mu_{0}^{q}\cdot\left(\frac{\lfloor s\rfloor}{n}-\frac{\mathbb{E}_{\mathbb{P}}[p_{1}(X^{n})^{2}]}{\varepsilon}\right)=(1-\varepsilon_{1})^{q}\mu_{0}^{q}\cdot\frac{\lfloor s\rfloor}{n}\left(1-\frac{o(1)}{\varepsilon}\right),$
which is $(1-o(1))(1-c(\varepsilon))\mu^{q-p}R^{p}$ with $c(\varepsilon)\to 0$ as $\varepsilon\to 0^{+}$, as claimed.
In the sequel we prove (48). Since
$\mathbb{E}_{\mathbb{P}}\left[(\theta_{1}-\mathbb{E}_{\mathbb{P}}[\theta_{1}|X^{n}])^{2}\right]=\mathbb{E}_{\mathbb{P}}[\theta_{1}^{2}]-\mathbb{E}_{\mathbb{P}}\left[\mathbb{E}_{\mathbb{P}}[\theta_{1}|X^{n}]^{2}\right]=\frac{\lfloor s\rfloor}{n}\mu_{0}^{2}-\mathbb{E}_{\mathbb{P}}\left[\mathbb{E}_{\mathbb{P}}[\theta_{1}|X^{n}]^{2}\right],$
it is equivalent to proving that $\mathbb{E}_{\mathbb{P}}\left[(\theta_{1}-\mathbb{E}_{\mathbb{P}}[\theta_{1}|X^{n}])^{2}\right]\geq(1-o(1))\frac{s\mu_{0}^{2}}{n}$. Let $\delta=\delta(\varepsilon)>0$ be any fixed constant such that
$(1+\delta)(1-\varepsilon)<1-\frac{\varepsilon}{2}.$ (49)
@@
$I(\theta;X^{\lambda})=\frac{1}{2}\int_{0}^{\lambda}\mathbb{E}\left[\|\theta-\mathbb{E}\left[\theta|X^{t}\right]\!\|_{2}^{2}\right]\mathrm{d}t=:\frac{1}{2}\int_{0}^{\lambda}\mathrm{mmse}(t)\mathrm{d}t.$
We denote by $I_{\mathbb{P}}(\theta;X^{\lambda})$ and $I_{\mathbb{Q}}(\theta;X^{\lambda})$ the mutual information under the permutation prior and the i.i.d. prior, respectively. Similarly we use the notations $(\mathrm{mmse}_{\mathbb{P}}(t),\mathrm{mmse}_{\mathbb{Q}}(t))$ for the minimum mean squared errors (MMSEs) under these priors. Note that our target is now equivalent to showing that $\mathrm{mmse}_{\mathbb{P}}(1)\geq(1-o(1))s\mu_{0}^{2}$.
By the I-MMSE formula,
For the difference in the mutual information, we establish the following inequality:
$0\leq I_{\mathbb{Q}}(\theta;X^{\lambda})-I_{\mathbb{P}}(\theta;X^{\lambda})\leq 10\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1}^{\lambda})^{\ell},$ (50)
where $I_{\chi^{2}}$ is the $\chi^{2}$ mutual information in Definition˜1.1. To show (50), dropping the superscript $\lambda$ for simplicity, we have
$I_{\mathbb{Q}}(\theta;X)-I_{\mathbb{P}}(\theta;X)$ $=nI_{\mathbb{Q}}(\theta_{1};X_{1})-\mathbb{E}_{\mathbb{P}_{\theta,X}}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}\mathbb{P}_{X_{i}|\theta_{i}}}{\mathrm{d}\mathbb{P}_{X}}\right]$
$=nI_{\mathbb{Q}}(\theta_{1};X_{1})-\mathbb{E}_{\mathbb{P}_{\theta,X}}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}\mathbb{P}_{X_{i}|\theta_{i}}}{\mathrm{d}\mathbb{Q}_{X}}\right]+{\sf KL}(\mathbb{P}_{X}\|\mathbb{Q}_{X})$
$=nI_{\mathbb{Q}}(\theta_{1};X_{1})-n\mathbb{E}_{\mathbb{P}_{\theta_{1},X_{1}}}\left[\log\frac{\mathrm{d}\mathbb{P}_{X_{1}|\theta_{1}}}{\mathrm{d}\mathbb{Q}_{X_{1}}}\right]+{\sf KL}(\mathbb{P}_{X}\|\mathbb{Q}_{X})$
$={\sf KL}(\mathbb{P}_{X}\|\mathbb{Q}_{X}),$
where the last step uses that $\mathbb{P}_{\theta_{1},X_{1}}=\mathbb{Q}_{\theta_{1},X_{1}}$. Since $\mathbb{P}_{X}$ and $\mathbb{Q}_{X}$ are the permutation mixture and the i.i.d. approximation, respectively, an intermediate step in the proof of the upper bound (4) (cf. Section˜4.2) shows that
$0\leq{\sf KL}(\mathbb{P}_{X}\|\mathbb{Q}_{X})\leq\chi^{2}(\mathbb{P}_{X}\|\mathbb{Q}_{X})\leq 10\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1})^{\ell},$
establishing the target inequality (50).
Based on (50), using the decreasing property of $t\mapsto\mathrm{mmse}(t)$ under both priors, we have
which rearranges to
The quantity $\mathrm{mmse}_{\mathbb{Q}}(1+\delta)$, this quantity corresponds to the Bayes risk under a rescaled i.i.d. prior $\left((1-\alpha)\delta_{0}+\alpha\delta_{\mu_{1}}\right)^{\otimes n}$ for the Gaussian mean vector, with
$\mu_{1}:=\sqrt{(1+\delta)(1-\varepsilon)}\mu=\sqrt{2(1+\delta)(1-\varepsilon)\log(1/R^{p})},\qquad\alpha:=\frac{\lfloor s\rfloor}{n}.$
As $\mu_{1}<\mu$ by (49) and $s=\omega(1)$, this Bayes risk has been characterized in [DJ94, Section 6] that
$\mathrm{mmse}_{\mathbb{Q}}(1+\delta)\geq\frac{(1-o(1))s\mu_{1}^{2}}{1+\delta}=(1-o(1))s\mu_{0}^{2}.$
Here the $(1+\delta)^{-1}$ factor accounts for discrepancy in estimating $\theta$ and $\sqrt{1+\delta}\cdot\theta$. Finally, since $\delta=\delta(\varepsilon)>0$ is a fixed constant and $s\mu_{0}^{2}=\omega(1)$, (51) proves the target claim $\mathrm{mmse}_{\mathbb{P}}(1)\geq(1-o(1))s\mu_{0}^{2}$ as long as we show that
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})=o(1).$ (52)
[…]
Consequently,
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})$ $\overset{\rm(a)}{=}\alpha(1-\alpha)\int\frac{(\mathrm{d}{\mathcal{N}}(\mu_{1},1)-\mathrm{d}{\mathcal{N}}(0,1))^{2}}{\mathrm{d}[\alpha{\mathcal{N}}(\mu_{1},1)+(1-\alpha){\mathcal{N}}(0,1)]}$
$\overset{\rm(b)}{=}\alpha(1-\alpha)\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{1}Z-\mu_{1}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{1}Z-\mu_{1}^{2}/2)+1-\alpha}\right],$
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{1}Z-\mu_{1}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{1}Z-\mu_{1}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{1}}\right\}}}\right]$
$\leq\frac{1}{1-\alpha}\left(1+\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(2\mu_{1}Z-\mu_{1}^{2}){\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{1}}\right\}}}\right]\right)$
$\overset{\rm(c)}{=}\frac{1}{1-\alpha}\left(1+e^{\mu_{1}^{2}}\cdot\mathbb{P}\left({\mathcal{N}}(2\mu_{1},1)\leq\left(1+\frac{\varepsilon}{4}\right)\mu_{1}\right)\right)$
$\overset{\rm(d)}{\leq}\frac{1}{1-\alpha}\left(1+\exp\left(\mu_{1}^{2}-\frac{1}{2}\left(1-\frac{\varepsilon}{4}\right)^{2}\mu_{1}^{2}\right)\right)=\frac{1}{1-\alpha}\left(1+\exp\left(\left(\frac{1}{2}+\frac{\varepsilon}{4}-\frac{\varepsilon^{2}}{32}\right)\mu_{1}^{2}\right)\right),$
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{1}Z-\mu_{1}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{1}Z-\mu_{1}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{1}}\right\}}}\right]$
$\leq\frac{1}{\alpha}\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(\mu_{1}Z-\mu_{1}^{2}/2){\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{1}}\right\}}}\right]$
$=\frac{1}{\alpha}\mathbb{P}\left({\mathcal{N}}(\mu_{1},1)>\left(1+\frac{\varepsilon}{4}\right)\mu_{1}\right)\leq\frac{1}{\alpha}\exp\left(-\frac{\varepsilon^{2}\mu_{1}^{2}}{32}\right).$
Since $\alpha\leq\exp\left(-\frac{(1+\varepsilon/2)\mu_{1}^{2}}{2}\right)$, we have
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})\leq\alpha+2\exp\left(-\frac{\varepsilon^{2}\mu_{1}^{2}}{32}\right)=o(1),$
as claimed. The proof of (52) is then complete.
C.6 Proof of Lemma˜2.2
It is clear that $\mathbb{P}_{n}$ and $\mathbb{P}_{n}^{\prime}$ are precisely the permutation mixtures in Theorem˜1.5, where the Laplace and Gaussian mechanisms correspond to the distribution class
${\mathcal{P}}_{1}=\left\{x+\mathrm{Lap}\left(\frac{1}{\varepsilon}\right):x\in[0,1]\right\},\quad{\mathcal{P}}_{2}=\left\{x+{\mathcal{N}}\left(0,\frac{1}{\varepsilon^{2}}\right):x\in[0,1]\right\},$
respectively. Straightforward computations yield
$\mathsf{D}_{\chi^{2}}({\mathcal{P}}_{1})=\frac{2e^{\varepsilon}+e^{-2\varepsilon}}{3}-1\leq\varepsilon^{2},\quad\text{for }\varepsilon\in[0,1].$
Since ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\vee(\Delta_{H^{2}}({\mathcal{P}})-1)\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$, Theorem˜1.5 yields
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})^{2}\leq\frac{3\varepsilon^{2}[e(1+\varepsilon^{2})]^{3\varepsilon^{2}}}{n}=O\left(\frac{\varepsilon^{2}}{n}\right),$
establishing the results for the Laplace mechanism. For the Gaussian mechanism, we similarly have
$\mathsf{D}_{\chi^{2}}({\mathcal{P}}_{2})=e^{\varepsilon^{2}}-1\leq 2\varepsilon^{2},\quad\text{for }\varepsilon\in[0,1],$
and the rest follows from Theorem˜1.5.
C.7 Proof of Lemma˜2.3
We first recall the alternative representations of $r^{\mathrm{S}}(\theta)$ and $r^{\mathrm{PI}}(\theta)$ in [GR09] (see also [Wei21]). Consider a “postulated Bayes model” where $\vartheta$ is a uniformly random permutation of a fixed vector $\theta$, and the observation $X\sim\otimes_{i=1}^{n}P_{\vartheta_{i}}$ conditioned on $\vartheta$. Due to the permutation-invariant structure of both ${\mathcal{D}}^{\mathrm{S}}$ and ${\mathcal{D}}^{\mathrm{PI}}$, it holds that
[…]
We next show that
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq\frac{M}{n}\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right].$ (53)
@@
$r^{\mathrm{S}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{S}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\Delta(\cdot)}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\Delta(X_{i}))\right]=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]\right],$
$r^{\mathrm{PI}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\widehat{\vartheta}_{i}(X))\right]=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right],$
@@
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ $\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\sup_{u}\left|\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]-\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right|\right]$
$\leq\frac{M}{n}\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right],$
@@
[…]
@@
[…]